OpenStudy (anonymous):

Express each of the following repeating decimals as a fraction: 0.426 (repeating sign over 6) as a fraction=? 0.426 (repeating sign over the 26) as a fraction=? 0.426 (repeating sign over 426) as a fraction=?

OpenStudy (anonymous):

\[0.42\overline{6}=\frac{42}{100} + \frac{0.\overline{6}}{100}=\frac{42}{100}+\frac{x}{100}\] \[x=0.\overline{6}\]\[10x=6.\overline{6}\]\[10x=6+x\]\[9x=6\]\[x=\frac{6}{9}\] \[0.42\overline{6}=\frac{42}{100}+\frac{\frac{6}{9}}{100}\]And then you can do some manipulation of fractions to make it a single fraction.

OpenStudy (anonymous):

Okay I see! and then would the next one be 10x=.26 ?\ Thank yoU!

OpenStudy (anonymous):

Yes, that's the jist of it. Then your fractions would be \[0.4\overline{26}=\frac{4}{10}+\frac{0.\overline{26}}{10}\], and you can replace the 0.26 with whatever fraction x is.

OpenStudy (anonymous):

Oh I see!! Thank you :)

OpenStudy (anonymous):

Oh actually, I misread your text a bit. You set \[x=0.\overline{26}\] first of all. Then your next step is to multiply by some power of 10, so that there is one block of the repeating digits before the decimal point, and then the repeating block after it, so:\[100x=26.\overline{26}\], and so on.

OpenStudy (anonymous):

oh i see now. thank you!